### 15 Mar 2026 v1

@@ Abstract
This article studies phonetic texture in Persian poetry as a literary-historical problem rather than as a by-product of meter or an auxiliary feature for classification. The corpus comprises 1,116,306 mesras from 31,988 poems and 83 poets drawn from a larger archive after restricting inference to five canonically named classical meters in the retained analytical subset and to well-supported poet-meter cells. Each mesra is represented as a grapheme-to-phoneme symbol stream and summarized through six measures: hardness, sonority, sibilance, vowel ratio, phoneme entropy, and consonant-cluster ratio.
@@
The article makes four contributions: it establishes a corpus-scale phonetic framework for Persian poetry; models poet differentiation under explicit formal controls while keeping meter legible as classical aruz; develops a multidimensional account of stylistic clustering and historical redistribution rather than a single hard-soft continuum; and demonstrates how computational phonetics can sharpen literary-historical interpretation without displacing it.
@@ 1 Introduction
This article asks four linked questions. First, do poet-level phonetic differences remain measurable after controlling for meter, poetic form, and line-length opportunity? Second, how are those differences organized in a multidimensional stylistic field rather than along a single hard-soft axis? Third, what century-by-century redistributions become visible when phonetic measures are read against Persian literary history? Fourth, how can those quantitative profiles be translated into historically credible readings of individual poets without collapsing literary value into sound alone?
The study answers those questions through a deliberately disciplined design: corpus restriction for comparability, symbolic phonetic representation at mesra scale, confound-aware statistical models, century interpretation, and case studies that return the measurements to literary history. The active inferential cohort comprises 1,116,306 mesras from 31,988 poems and 83 poets. Ten of the main case-study poets fall squarely inside that balanced cohort; Ferdowsi is added as a high-support epic reference; Shahriar and Khayyam are retained as descriptive extensions whose historical importance exceeds their fit with the strongest inferential frame. This architecture lets the article compare poets rigorously while remaining explicit about evidential boundaries. The article makes four contributions:
@@ 2 Related Work
Research on Persian poetry in digital humanities is now rich enough that the main question is no longer whether large-scale comparison is possible, but what kind of literary object each computational design actually makes visible. The present study belongs at the intersection of corpus infrastructure, stylometry and prosody, phonetic representation, and sound-centered digital humanities, but it addresses a gap that those strands have not yet resolved together.
@@ 2.1 Persian Corpora and Literary Modeling
Corpus construction has been the enabling condition of recent Persian DH. Early work by Asgari and Chappelier showed that Persian poetic corpora could sustain large-scale analysis of literary tendencies rather than only isolated text-processing tasks [1]. That infrastructural turn was extended decisively by Raji et al., whose large corpus and metadata architecture made poet-, genre-, and period-level comparison practically viable [2]. More recent systems such as PARSI and NAZM have moved from corpus assembly to literary profiling, showing that Persian poets and traditions can be compared through stylometric, semantic, and metrical features [3, 4]. Work on thematic similarity and era-sensitive classification likewise demonstrates that Persian poetic corpora can sustain temporal and stylistic discrimination at poem scale [5, 6]. This literature establishes the feasibility of poet-level modeling, but its primary objects are usually authorship, chronology, theme, or broad style rather than phonetic surface as a humanities object in its own right.
@@ 2.2 Prosody, Stylometry, and Authorship
The closest methodological neighbors to this article come from stylometry, authorship studies, computational prosody, and phonetic representation. Quantitative work on attribution, such as Ishikawa’s study of the Attar corpus, demonstrates how statistical evidence can enter philological debate without claiming to replace it [7]. Prosodic studies likewise show that Persian aruz is computationally tractable, as in Shahrestani and Chehreghani’s recognition work [8]. At the level of phonetic reconstruction, Qharabagh et al. demonstrate that Persian grapheme-to-phoneme conversion is itself an active computational problem rather than a transparent preprocessing step [9]. Parallel work in Arabic and multilingual meter modeling reinforces the same point: formal structure can be learned computationally and matters deeply for literary analysis [10, 11].
@@ 2.3 Sound-Centered Digital Humanities
Outside Persian studies, sound-centered digital humanities has provided both conceptual orientation and methodological precedent. Clement’s call to “sound for meaning” argued that aural patterning belongs to interpretation rather than to a merely decorative supplement [12]. Poemage turned that principle into an exploratory visual environment for sonic structure within poems [13]. More recently, Meng et al. showed that phonetic vectors can distinguish major Chinese poets under strong formal constraint, demonstrating that sound can support historically specific comparison instead of only generic euphony claims [14].
@@ 2.4 Gap Addressed Here
The gap addressed by this article is therefore specific. Persian DH already has strong work on corpus infrastructure, stylometry, prosody, phonetic engineering, generative modeling, and broader AI-oriented literary exploration [15, 16, 17, 18, 9]. Sound-centered DH already shows how phonetic patterning can matter for literary interpretation. What is missing is a Persian framework that treats phonetic texture itself as the object of explanation, while controlling explicitly for meter, form, and line-length opportunity and then returning the results to literary history.
This study addresses that gap by shifting the inferential target. Sound is not used primarily as an auxiliary predictor for classification. Instead, it is modeled as a multidimensional field in which poets, meters, and centuries occupy historically intelligible positions. Its novelty lies not in rejecting prior Persian DH, but in assembling its strongest strands into a single argument: phonetic representation, formal control, meter interpretation, stylistic clustering, and literary-historical reading belong in the same analytical frame if we want to hear the corpus historically rather than merely process it computationally.
@@ 3 Data and Preprocessing
The article begins from a large mesra-level archive, but it does not treat the archive as one homogeneous evidential field. In Persian poetry, phonetic comparison is inseparable from meter, poetic form, and line length; preprocessing is therefore part of the scholarly design, not merely a technical prelude. The aim is to define a domain in which poet-level comparison remains historically legible and formally controlled.
@@ 3.1 Corpus Scope and Inferential Domain
The inferential cohort is further limited to verse in five canonically named classical meters of Persian aruz: mutaqarib, hazaj, ramal, mujtass, and muzari’. In Persian prosody, a bahr is a line-level quantitative pattern of long (-) and short ($\cup$) syllables distributed across the mesra; individual realizations admit historically recognized sub-patterns and zehafat rather than free variation. [^1] For the canonical nomenclature of Persian quantitative prosody, including ramal, hazaj, rajaz, mutaqarib, muzari’, mujtass, khafif, munsarih, and muqtadab, see [19, 20, 21]. Spellings such as mutagharib for mutaqarib, modar for muzari’, mujtath for mujtass, and maqtadab for muqtadab reflect transliteration choice rather than distinct meters. In this article the subsequent five-way grouping is a post-hoc analytic convenience: the canonical literary category is the named bahr, while corpus codes such as M01–M05 identify specific retained sub-patterns. These are not arbitrary computational bins. They are historically documented meters that dominate the metrically annotated archive and form the part of the corpus in which controlled comparison remains statistically credible and literarily legible.
@@
The final cohort retains 38.6% of broad-corpus mesras and 83 of 216 poets. Read from the metrically annotated subset rather than the full archive, the five retained meters account for 77.6% of all mesras with usable meter/form metadata, and after the support threshold they constitute the entire inferential cohort. Their shares inside the final cohort are 30.3% for mutaqarib, 21.1% for hazaj, 17.8% for ramal, 17.3% for mujtass, and 13.5% for muzari’. The gain is not statistical tidiness alone. It is literary discipline: poets are compared inside shared formal conditions, and later interpretation is not driven by marginal meter cells, unbalanced century tails, or incomplete metadata.
@@ 3.3 Corpus Structure
Figure 1 now summarizes the analytical population through two coordinated views. A figure header reports the scale of the balanced cohort directly, avoiding the distracting indexed-count panel used in earlier drafts. Panel A turns to prosody by showing that the comparison set is concentrated in the five retained classical meters. Panel B then shows the historical coverage of the metadata-linked subset, which is strongest from the tenth through seventeenth centuries and preserves a smaller but important nineteenth-century tail.
Figure 1: Structure of the analytical corpus. The header reports the scale of the balanced cohort directly: 1,116,306 mesras, 31,988 poems, and 83 poets across the five retained classical meters. In mesra terms these meters account for 77.6% of rows with usable meter/form metadata and 100% of the final inferential cohort. Panel A shows poem counts and shares across mutaqarib, hazaj, ramal, mujtass, and muzari’, identifying the prosodic field within which controlled comparison is possible. Panel B plots poems per century for the metadata-linked subset, showing the historical span on which the century analysis rests.
Taken together, these distributions clarify the interpretive stance of the article. The models are not estimating “Persian poetry” in the abstract. They are estimating phonetic differentiation inside a historically broad but formally disciplined subset dominated by a small number of aruz repertoires. That is a narrower claim than canon-wide representation, but a much stronger claim than unrestricted corpus description.
@@ 3.4 Interpretive Consequence
The preprocessing argument is therefore straightforward. A large corpus becomes a persuasive humanities object only when the boundaries of comparison are as explicit as the scale of the archive. By restricting the inferential domain, the article gives up some breadth in order to gain historical legibility: epic diction is compared with lyric diction inside identifiable meters, later poets are not allowed to dominate thinly supported cells, and century claims remain tethered to known coverage. That trade-off is essential to everything that follows.
@@ 4.1 Symbol Streams and Boundary Rules
Each mesra is converted into an ordered sequence of atomic symbols after minimal cleanup. Word boundaries are preserved, mesra boundaries are split only on explicit structural delimiters, and in-token phonological symbols are not mistaken for separators. This rule is essential in Persian G2P material, where some characters can function as phonological units in one context and as boundary markers in another. The representation is designed for transparency rather than compression: later claims about elevated sibilance or reduced sonority remain traceable to readable segmental distributions.
@@ 4.2 Feature Mapping
The weighting scheme matters interpretively. A poet can become “harder” either by accumulating more obstruent pressure overall or by favoring specifically dorsal and uvular segments; conversely, sonority can rise through vowels and glides without eliminating all consonantal edge. The paper therefore treats the representation as an interpretable feature map rather than as an opaque embedding.
@@ 4.4 Representational Limit
The representational claim of the article is precise. Hardness, sonority, sibilance, and related measures describe encoded phonetic texture in the corpus representation, not direct recordings of historical recitation. This limit is important, but it does not weaken the study. It clarifies the inferential object. Once line boundaries, named meters, poet identities, and phonetic symbols are stabilized, corpus statistics can be returned to literary history without pretending to measure performance itself.
@@ 5 Metrics and Statistical Models
The analysis models Persian poetic sound through six complementary measures defined over the mesra-level symbol stream. The point is not to collapse style into a single acoustic index, but to preserve several dimensions that criticism often gathers under broader labels such as firmness, fluency, brilliance, density, or softness.
@@ 5.4 Interpretive Hierarchy
The paper maintains a clear evidential hierarchy. Hardness, sonority, and sibilance support the strongest inferential claims because they are modeled directly. Vowel ratio, entropy, and cluster ratio mainly clarify how poet profiles are assembled. Principal-component projection and fingerprint plots are therefore descriptive aids, not independent inferential results. They matter because they make the geometry of style readable, but they are always interpreted against the controlled models, the meter profiles, and the historical argument.
That hierarchy is essential for responsible reading. Statistical significance alone is not treated as a literary claim. When the paper moves from coefficients to interpretation, it does so through effect magnitude, cross-metric coherence, meter context, and historical plausibility. The method is designed to discipline literary judgment, not to mechanize it.
@@ 6 Results
The results move from controlled differentiation to historical interpretation. The first layer establishes that the core phonetic measures are well behaved and that poet effects persist after formal control. The second reads century trends against the actual temporal coverage of the corpus: the tenth through seventeenth centuries, with a smaller nineteenth-century return. The third places poets in a multidimensional stylistic space, and the fourth returns meter from control variable to prosodic institution.
@@ 6.2 Century-by-Century Historical Redistribution
The nineteenth century appears in the corpus as a smaller but revealing tail. Its higher sibilance and renewed sonority fit a period in which revivalist rhetoric, didactic-public address, devotional circulation, and expanding print culture redistribute older resources rather than simply repeating them. Safi Alishah, Ashofteh Shirazi, and Bolandeghbal exemplify three distinct nineteenth-century solutions: exhortative mystical contour, calibrated bazgasht modulation, and Sufi-didactic public bite. The century curves therefore track changes in literary ecology, performance context, and poetic register, not a teleological sonic destiny.
@@ 6.3 Phonetic Stylistic Space and Cluster Logic
Figure 6: Phonetic stylistic space of Persian poets. Grey points show the wider field; colored points mark the focal case studies and historical anchors. The cluster legend is placed outside the plotting area so the labeled cases remain readable.
@@
Table 2 summarizes these zones as descriptive archetypes. The point is comparative rather than taxonomic. The table shows that lyric openness, rhetorical force, sibilant brightness, and experimental density recur across traditions, but they do so through historically distinct combinations of the six metrics.
@@ 6.4 Meters as Prosodic Institutions
Meter is not only a control variable in this article. It is one of the principal institutions through which sonic expectation is organized. Table 3 translates the internal meter codes into named classical aruz meters and the exact retained sub-patterns, and reports poem-level phonetic means for the five dominant meters in the balanced subset. [^2] In the analytic coding used here, M01–M05 correspond to the sub-patterns mutaqarib-i musamman-i mahzuf, hazaj-i musaddas-i mahzuf, ramal-i musamman-i mahzuf, mujtass-i musamman-i makhbun-i mahzuf, and muzari’-i musamman-i akhrab-i makfuf-i mahzuf. Following [19, 20, 21], these are best understood as specific sub-patterns or permissible variants within named bahrs, not as members of an abstract literary “family.” The five meters were selected because they dominate the metrically annotated archive and therefore support both statistical generalization and historically intelligible interpretation: together they account for 77.6% of mesras with usable meter/form metadata and 100% of the final support-filtered cohort.
@@
Classical prosody and literary history make the same selection intelligible. Mutaqarib is the canonical epic meter of the Shahnameh tradition and remains strongly associated with heroic or declarative narrative; hazaj is prominent in romance, chantable lyric extension, and later devotional reuse; ramal is central to ghazal and mystical lyric from the high-classical period onward; mujtass regularly supports compact ghazal brightness; and muzari’ is especially visible in rhetorically wrought late-classical, Indo-Persian, and revivalist verse. [^3] For the canonical metrical inventory and the status of these forms as named Persian bahrs, see [19, 20, 21]. The genre and period associations emphasized here synthesize those standard descriptions with the distributions observed in this corpus. The article therefore uses historically documented meters rather than approximate metrical classes, while still recognizing that the five-way grouping is an analytic restriction imposed for inferential stability.
@@
The within-meter evidence supports the same interpretation. Mean pairwise correlation of poet effects is 0.58 for hardness, 0.68 for sonority, and 0.31 for sibilance. Sonority is therefore the most portable dimension across metrical envelopes, hardness remains moderately portable, and sibilance is the most meter-sensitive. That pattern is historically plausible. Vocalic openness can travel across several repertoires, whereas strident contour tends to sharpen most strongly when a meter already licenses it. The full within-meter standardized-effects tables are therefore moved to Appendix B, where readers can inspect the complete coefficient matrices for hardness, sonority, and sibilance without interrupting the main historical argument.

### 01 Apr 2026 v2

@@ Abstract
This study examines phonetic texture in Persian poetry as a literary-historical problem rather than as a by-product of meter or an auxiliary feature for classification. The corpus comprises 1,116,306 mesras from 31,988 poems and 83 poets drawn from a larger archive after restricting inference to five canonically named classical meters in the retained analytical subset and to well-supported poet-meter cells. Each mesra is represented as a grapheme-to-phoneme symbol stream and summarized through six measures: hardness, sonority, sibilance, vowel ratio, phoneme entropy, and consonant-cluster ratio.
@@
Taken together, the results provide a corpus-scale phonetic framework for Persian poetry. They estimate poet differentiation under explicit formal controls while keeping meter legible as classical aruz, model stylistic clustering and historical redistribution beyond a single hard-soft continuum, and show how computational phonetics can sharpen literary-historical interpretation without displacing it.
@@ 1 Introduction
The central questions are whether poet-level phonetic differences remain measurable after controlling for meter, poetic form, and line-length opportunity; how those differences are organized in a multidimensional stylistic field rather than along a single hard-soft axis; what century-by-century redistributions become visible when phonetic measures are read against Persian literary history; and how such quantitative profiles can be translated into historically credible readings of individual poets without collapsing literary value into sound alone.
To address them, the analysis combines corpus restriction for comparability, symbolic phonetic representation at mesra scale, confound-aware statistical models, century interpretation, and case studies that return the measurements to literary history. The active inferential cohort comprises 1,116,306 mesras from 31,988 poems and 83 poets. Ten of the main case-study poets fall inside that balanced cohort; Ferdowsi serves as a high-support epic reference; Shahriar and Khayyam are retained as descriptive extensions because their historical importance exceeds their fit with the strongest inferential frame. This design allows rigorous comparison while keeping evidential boundaries explicit. Four features are central:
@@ 2 Related Work
Research on Persian poetry in digital humanities is now rich enough that the main question is no longer whether large-scale comparison is possible, but what kind of literary object each computational design actually produces. The present study belongs at the intersection of corpus infrastructure, stylometry and prosody, phonetic representation, and sound-centered digital humanities, but it addresses a gap that those strands have not yet resolved together.
@@ 2.1 Persian Corpora and Literary Modeling
Corpus construction has been the enabling condition of recent Persian DH. Early work by Asgari and Chappelier showed that Persian poetic corpora could sustain large-scale analysis of literary tendencies rather than only isolated text-processing tasks [asgari2013linguistic]. That infrastructural turn was extended decisively by Raji et al., whose large corpus and metadata architecture made poet-, genre-, and period-level comparison practically viable [raji2023corpus]. More recent systems such as PARSI and NAZM have moved from corpus assembly to literary profiling, showing that Persian poets and traditions can be compared through stylometric, semantic, and metrical features [shahnazari2025parsi, shahnazari2025nazm]. Work on thematic similarity and era-sensitive classification likewise demonstrates that Persian poetic corpora can sustain temporal and stylistic discrimination at poem scale [akef2020doc2vec, ruma2022hafez]. This literature establishes the feasibility of poet-level modeling, but its primary objects are usually authorship, chronology, theme, or broad style rather than phonetic surface as a humanities object in its own right.
@@ 2.2 Prosody, Stylometry, and Authorship
The closest methodological neighbors to this study come from stylometry, authorship studies, computational prosody, and phonetic representation. Quantitative work on attribution, such as Ishikawa’s study of the Attar corpus, demonstrates how statistical evidence can enter philological debate without claiming to replace it [ishikawa2025attar]. Prosodic studies likewise show that Persian aruz is computationally tractable, as in Shahrestani and Chehreghani’s recognition work [shahrestani2025prosody]. At the level of phonetic reconstruction, Qharabagh et al. demonstrate that Persian grapheme-to-phoneme conversion is itself an active computational problem rather than a transparent preprocessing step [qharabagh2025llm]. Parallel work in Arabic and multilingual meter modeling reinforces the same point: formal structure can be learned computationally and matters deeply for literary analysis [delarosa2021transformers, qarah2024arapoembert].
@@ 2.3 Sound-Centered Digital Humanities
Outside Persian studies, sound-centered digital humanities has provided both conceptual orientation and methodological precedent. Clement’s call to “sound for meaning” argued that aural patterning belongs to interpretation rather than to a merely decorative supplement [clementSoundingMeaning]. Poemage turned that principle into an exploratory visual environment for sonic structure within poems [mccurdy2016poemage]. More recently, Meng et al. showed that phonetic vectors can distinguish major Chinese poets under strong formal constraint, demonstrating that sound can support historically specific comparison instead of only generic euphony claims [meng2025listening].
@@ 2.4 Gap Addressed Here
The gap is specific. Persian DH already has strong work on corpus infrastructure, stylometry, prosody, phonetic engineering, generative modeling, and broader AI-oriented literary exploration [khanmohammadi2021prose2poem, panahandeh2023tppoet, moghadam2018newpoet, rastimeymandi2024opportunities, qharabagh2025llm]. Sound-centered DH already shows how phonetic patterning can matter for literary interpretation. What is missing is a Persian framework that treats phonetic texture itself as the object of explanation, while controlling explicitly for meter, form, and line-length opportunity and then returning the results to literary history.
Accordingly, sound is not used primarily as an auxiliary predictor for classification. It is modeled instead as a multidimensional field in which poets, meters, and centuries occupy historically intelligible positions. The novelty lies not in rejecting prior Persian DH, but in assembling its strongest strands into a single argument: phonetic representation, formal control, meter interpretation, stylistic clustering, and literary-historical reading belong in the same analytical frame if the corpus is to be heard historically rather than merely processed computationally.
@@ 3 Data and Preprocessing
The analysis begins from a large mesra-level archive, but it does not treat the archive as one homogeneous evidential field. In Persian poetry, phonetic comparison is inseparable from meter, poetic form, and line length; preprocessing is therefore part of the scholarly design, not merely a technical prelude. The aim is to define a domain in which poet-level comparison remains historically legible and formally controlled.
@@ 3.1 Corpus Scope and Inferential Domain
The inferential cohort is further limited to verse in five canonically named classical meters of Persian aruz: mutaqarib, hazaj, ramal, mujtass, and muzari’. In Persian prosody, a bahr is a line-level quantitative pattern of long (-) and short ($\cup$) syllables distributed across the mesra; individual realizations admit historically recognized sub-patterns and zehafat rather than free variation. [^1] For the canonical nomenclature of Persian quantitative prosody, including ramal, hazaj, rajaz, mutaqarib, muzari’, mujtass, khafif, munsarih, and muqtadab, see [iranicaAruz, elwellSutton1976persianmetres, britannicaPersianClassicalPoetry]. Spellings such as mutagharib for mutaqarib, modar for muzari’, mujtath for mujtass, and maqtadab for muqtadab reflect transliteration choice rather than distinct meters. Here the subsequent five-way grouping is a post-hoc analytic convenience: the canonical literary category is the named bahr, while corpus codes such as M01–M05 identify specific retained sub-patterns. These are not arbitrary computational bins. They are historically documented meters that dominate the metrically annotated archive and form the part of the corpus in which controlled comparison remains statistically credible and literarily legible.
@@
The final cohort retains 38.6% of broad-corpus mesras and 83 of 216 poets. Read from the metrically annotated subset rather than the full archive, the five retained meters account for 77.6% of all mesras with usable meter/form metadata, and after the support threshold they constitute the entire inferential cohort. Their shares inside the final cohort are 30.3% for mutaqarib, 21.1% for hazaj, 17.8% for ramal, 17.3% for mujtass, and 13.5% for muzari’. This restriction improves comparability rather than simply tidying the data: poets are compared under shared formal conditions, and later interpretation is less exposed to marginal meter cells, thin century tails, or incomplete metadata.
@@ 3.3 Corpus Structure
Figure 1 presents the analytical population in two views. The header reports the size of the balanced cohort directly. Panel A shows how poems are distributed across the five retained classical meters. Panel B shows the century distribution of the metadata-linked subset, which is densest from the tenth through seventeenth centuries and still includes a smaller nineteenth-century tail.
Figure 1: Structure of the analytical corpus. The header reports the scale of the balanced cohort directly: 1,116,306 mesras, 31,988 poems, and 83 poets across the five retained classical meters. In mesra terms these meters account for 77.6% of rows with usable meter/form metadata and 100% of the final inferential cohort. Panel A shows poem counts and shares across mutaqarib, hazaj, ramal, mujtass, and muzari’, defining the prosodic range of the controlled comparison. Panel B plots poems per century for the metadata-linked subset, showing the historical span covered by the century analysis.
These distributions show what the models do and do not represent. They estimate phonetic differentiation within a historically broad but formally restricted subset dominated by a small number of aruz repertoires. That is narrower than a claim about the canon as a whole, but more defensible than an unrestricted corpus description.
@@ 3.4 Interpretive Consequence
The inferential domain is restricted so that the boundaries of comparison remain explicit. That choice gives up some breadth, but it makes the comparison historically legible: epic diction is compared with lyric diction within identifiable meters, thin poet-meter cells do not drive the results, and century claims remain tied to observed coverage. That trade-off is essential to everything that follows.
@@ 4.1 Symbol Streams and Boundary Rules
Each mesra is converted into an ordered sequence of atomic symbols after minimal cleanup. Word boundaries are preserved, mesra boundaries are split only on explicit structural delimiters, and in-token phonological symbols are not mistaken for separators. This rule is essential in Persian G2P material, where some characters can function as phonological units in one context and as boundary markers in another. The representation is designed for transparency rather than compression: later claims about elevated sibilance or reduced sonority remain traceable to explicit segmental distributions.
@@ 4.2 Feature Mapping
The weighting scheme matters interpretively. A poet can become “harder” either by accumulating more obstruent pressure overall or by favoring specifically dorsal and uvular segments; conversely, sonority can rise through vowels and glides without eliminating all consonantal edge. The representation is therefore treated as an interpretable feature map rather than as an opaque embedding.
@@ 4.4 Representational Limit
The representational claim is narrow but important. Hardness, sonority, sibilance, and related measures describe encoded phonetic texture in the corpus representation, not direct recordings of historical recitation. This limit clarifies the inferential object. Once line boundaries, named meters, poet identities, and phonetic symbols are stabilized, corpus statistics can be returned to literary history without pretending to measure performance itself.
@@ 5 Metrics and Statistical Models
The analysis models Persian poetic sound through six complementary measures defined over the mesra-level symbol stream. Rather than collapsing style into a single acoustic index, it preserves several dimensions that criticism often gathers under broader labels such as firmness, fluency, brilliance, density, or softness.
@@ 5.4 Interpretive Hierarchy
The evidential hierarchy is clear. Hardness, sonority, and sibilance support the strongest inferential claims because they are modeled directly. Vowel ratio, entropy, and cluster ratio mainly clarify how poet profiles are assembled. Principal-component projection and fingerprint plots are descriptive aids, not independent inferential results. They clarify the geometry of style, but they are always interpreted against the controlled models, the meter profiles, and the historical argument.
That hierarchy is essential for responsible reading. Statistical significance alone is not treated as a literary claim. When the analysis moves from coefficients to interpretation, it does so through effect magnitude, cross-metric coherence, meter context, and historical plausibility. The method is designed to discipline literary judgment, not to mechanize it.
@@ 6 Results
The results proceed from controlled contrasts to century-level redistribution, stylistic-space projection, and meter profiles.
@@ 6.2 Century-by-Century Historical Redistribution
The nineteenth century appears in the corpus as a smaller but still informative tail. Its higher sibilance and renewed sonority fit a period in which revivalist rhetoric, didactic-public address, devotional circulation, and expanding print culture redistribute older resources rather than simply repeating them. Safi Alishah, Ashofteh Shirazi, and Bolandeghbal exemplify three distinct nineteenth-century solutions: exhortative mystical contour, calibrated bazgasht modulation, and Sufi-didactic sharpness. The century curves therefore track changes in literary ecology, performance context, and poetic register rather than a single linear sonic trajectory.
@@ 6.3 Phonetic Stylistic Space and Cluster Logic
Figure 6: Phonetic stylistic space of Persian poets. Grey points show the wider field; colored points mark the focal case studies and historical anchors, with the cluster legend placed outside the plotting area.
@@
Table 2 summarizes these zones as descriptive archetypes. It is meant as a comparative aid rather than a taxonomy. The table shows that lyric openness, rhetorical force, sibilant brightness, and experimental density recur across traditions, but they do so through historically distinct combinations of the six metrics.
@@ 6.4 Meters as Prosodic Institutions
Meter is not only a control variable here. It is one of the principal institutions through which sonic expectation is organized. Table 3 translates the internal meter codes into named classical aruz meters and the exact retained sub-patterns, and reports poem-level phonetic means for the five dominant meters in the balanced subset. [^2] In the analytic coding used here, M01–M05 correspond to the sub-patterns mutaqarib-i musamman-i mahzuf, hazaj-i musaddas-i mahzuf, ramal-i musamman-i mahzuf, mujtass-i musamman-i makhbun-i mahzuf, and muzari’-i musamman-i akhrab-i makfuf-i mahzuf. Following [iranicaAruz, elwellSutton1976persianmetres, britannicaPersianClassicalPoetry], these are best understood as specific sub-patterns or permissible variants within named bahrs, not as members of an abstract literary “family.” The five meters were selected because they dominate the metrically annotated archive and therefore support both statistical generalization and historically intelligible interpretation: together they account for 77.6% of mesras with usable meter/form metadata and 100% of the final support-filtered cohort.
@@
Classical prosody and literary history make the same selection intelligible. Mutaqarib is the canonical epic meter of the Shahnameh tradition and remains strongly associated with heroic or declarative narrative; hazaj is prominent in romance, chantable lyric extension, and later devotional reuse; ramal is central to ghazal and mystical lyric from the high-classical period onward; mujtass regularly supports compact ghazal brightness; and muzari’ is especially visible in rhetorically wrought late-classical, Indo-Persian, and revivalist verse. [^3] For the canonical metrical inventory and the status of these forms as named Persian bahrs, see [iranicaAruz, elwellSutton1976persianmetres, britannicaPersianClassicalPoetry]. The genre and period associations emphasized here synthesize those standard descriptions with the distributions observed in this corpus. Historically documented meters are therefore used rather than approximate metrical classes, while still recognizing that the five-way grouping is an analytic restriction imposed for inferential stability.
@@
The within-meter evidence supports the same interpretation. Mean pairwise correlation of poet effects is 0.58 for hardness, 0.68 for sonority, and 0.31 for sibilance. Sonority is therefore the most portable dimension across metrical envelopes, hardness remains moderately portable, and sibilance is the most meter-sensitive. That pattern is historically plausible. Vocalic openness can travel across several repertoires, whereas strident contour tends to sharpen most strongly when a meter already licenses it. The full within-meter standardized-effects tables appear in Appendix B.